Please need help!!! Thanks a lot
The first four questions refer to the circuit show
Answer #1

20-102 2 = 2XL z = √ 12 + x - x 5 = 2XL = √ R²+(XL-XL) 2 e a IR = I (UTS) XL Vize Vis Vimax = Clow) = 19.5V IWL lov =2 Viman

INR - Imare R - 19.50 0.0316 - R-ve Iman 616.68 o- tai 6325(0.05) - 6328 (1 ) 616.68 - 14.5 Vmax max = - V22+(WL wel I max 2